Test the series

Σ(-1)^(n-1)× sinnx/n√n for absolute and

n=1 conditional convergence


1
Expert's answer
2021-05-06T16:16:09-0400

Let us test the series "\sum_{n=1}^{\infty}\frac{(-1)^{n-1}\sin nx}{n\sqrt{n}}" for absolute and conditional convergence.


Since for any "x\in\mathbb R" we have that "|\frac{(-1)^{n-1}\sin nx}{n\sqrt{n}} |\le \frac{1}{n\sqrt{n}}=\frac{1}{n^{\frac{3}{2}}}" and the s-series "\sum_{n=1}^{\infty}\frac{1}{n^{\frac{3}{2}}}" is convergenent for "s=\frac{3}{2}>1", we conclude that by comparison test the series "\sum_{n=1}^{\infty}\frac{(-1)^{n-1}\sin nx}{n\sqrt{n}}" is absolute convergent for any "x\in\mathbb R", and hence the series is also conditional convergent for any "x\in\mathbb R".